Light travels from air into an optical fiber with an index of refraction of 1.44. 

(a) In which direction does the light bend? (b) If the angle of incidence on the

end of the fiber is 22 o , what is the angle of refraction inside the fiber? (c) 

Sketch the path of light as it changes media.


1
Expert's answer
2022-06-07T17:34:21-0400

"\frac{\sin\alpha}{\sin\beta}=n,"

"\beta=15°."